A golf ball of mass 0.045 kg is hit off the tee at a speed of 36 m/s . The golf club was in contact with the ball for 3.50×10−3 s .Find the average force exerted on the ball by the golf club.
Answer #1

Data Given

Mass m = 0.045 kg, velocity v = 36 m/s, time t = 3.50 X 10-3 s

First of all lets calculate the impulse on the ball

J = \Delta p = m \Delta v = 0.045 kg (36 m/s - 0 m/s )= 1.62 N.s

We also know that

F_{Avg} = \frac{J}{\Delta t}

F_{Avg} = \frac{ 1.62 N.s}{3.50\times 10^{-3}s}= 462.86 N